Jennifer ivested three times as much in an account that pays 5% simple interest as she did in an account that pays 2%. If her total annual interest is $35.70, how much did she invest in each account?
Answer by checkley77(12844)

.05*3x+.02x=35.70
.15x+.02x=35.70
.17x=35.70
x=35.70/.17
x=$210 amount invested @ 2%
3*210=630 amount invested @ 5%.
Proof:
.05*630+.02*210=35.70
31.50+4.2=35.70
35.70=35.70